plzzzzzzzzzz help me
In the cafeteria of Karachi University 40% drink Coca-cola,50% drink pepsi-cola and 25% drink both. If a person is selected ,find the probability he/she drinks either coca-cola or pepsi-cola.

Answer by edjones(8007)

The 25% who drink both are included in the 40% and 50%.
Since this is an either-or question we must eliminate th 25% from the statistics before adding.
40-25=15%
50-25=25%
15+25=40% drink either coke or pepsi (but not both).
.
Ed